OpenStudy (anonymous):

m angle 1 = (4x+12) degrees, m angle 2 = (7x+3) degrees, and m angle 3 = 6y degrees, then x = ?, m angle 1 = ?, m angle 2= ?, and y =?

OpenStudy (anonymous):

are these interior angles of a tirangle?

OpenStudy (anonymous):

im doing stuff with linear and vertical angles

OpenStudy (anonymous):

picture?

OpenStudy (anonymous):

Is this in a triangle?

OpenStudy (anonymous):

 do you see how angle 1 and angle 2, when added = 180 degrees?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then we can solve for x, (4x+12) + (7x+3) = 180

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (anonymous):

add it all? and get 26=x?

OpenStudy (anonymous):

close... (4x+12) + (7x+3) = 180 becomes 4x+12+7x+3=180 11x+12+3=180 11x + 15 = 180 11x = 180 - 15 11x = 165 x = ...

OpenStudy (anonymous):

how does the 15 go over to the other side and subtract from 180?

OpenStudy (anonymous):

great question, its a method/rule used to 'isolate x' 11x + 15 = 180 11x + 15 - 15 = 180 -15 11x = 165 If I subtract any number from one side, i have to do it to the other side. make sense?

OpenStudy (anonymous):

yes, so does x=11 then?

OpenStudy (anonymous):

close... so similar to how I subtracted 15 from both sides, i'm going to divide by 11 on both sides 11x = 165 (11x)/11 = 165/11 [the 11's on the left cancel out] x = 165/11

OpenStudy (anonymous):

so x=15.

OpenStudy (anonymous):

but how would i find what m angle 1, m angle 2 and y equal?

OpenStudy (anonymous):



OpenStudy (anonymous):

but how would i find y? m angle 3=6y